TheSmartOne (thesmartone):

Solve the equation after making an appropriate substitution. \(\sqrt{p}-2\sqrt[4]{p}-8=0\)

OpenStudy (kirbykirby):

Oh sorry .. actually, let \(y= \sqrt[4]{p}=p^{1/4}\) Then you have \(y^2-2y-8=0\)

TheSmartOne (thesmartone):

\(\sqrt{p}-2\sqrt[4]{p}-8=0\) was the question.